# Tunable ultrastrong magnon–magnon coupling and spin Hall magnetoresistance in a van der Waals antiferromagnet

### v3

@@
Tunable ultrastrong magnon–magnon coupling and spin Hall magnetoresistance in a van der Waals antiferromagnet
C. W. F. Freeman1,2,∗, Z. Xue1, H. Youel1, A.K. Budniak3,4, H. De Libero5, T. Thomson5,
@@
Antiferromagnetic (AFM) magnons and the manipulation of magnetisation via spin currents in van der Waals (vdW) materials offer substantial potential for applications in magnonics and spintronics. In this study, we demonstrate ultrastrong magnon-magnon coupling in the GHz regime within a vdW AFM, achieving a maximum normalised coupling strength of 0.89. Our investigation shows the tunability of coupling strength through temperature-dependent magnetic anisotropies. We compare coupling strength values derived from gap size with those calculated directly through the coupling parameter, showing that gap size as a measure of coupling strength is limited when intrinsic symmetry breaking induced magnon mode coupling is present. Additionally, analytical calculations show the possibility to reach the deep-strong coupling regime, with normalised coupling strengths greater than 1 by tuning the anisotropy values. Finally, we report on spin Hall magnetoresistance in a vdW AFM-based device and study the field and temperature dependence. These findings highlight the transformative potential of vdW AFMs in advancing the field of spin-based technologies.
@@
Keywords: Spintronics, Magnonics, 2D Materials, Antiferromagnets, Spin Orbitronics
@@
Within the expansive range of materials explored for spintronic applications, antiferromagnets (AFMs) have emerged as compelling candidates owing to their unique properties and potential to address challenges associated with traditional ferromagnetic (FM) devices [1, 2]. Characterized by the antiparallel alignment of neighboring magnetic moments, net zero magnetisation and atomistic exchange coupling strength, AFMs offer inherent advantages such as insensitivity to external magnetic fields and ultrafast spin dynamics [3]. However, the intricate nature of spin dynamics in AFMs presents both opportunities and challenges, underscoring the need for further investigation to fully harness their potential in spintronics.
@@
Van der Waals (vdW) materials are layered materials that can be exfoliated down to monolayer structures relatively easily under controlled experimental conditions [4]. The recent discovery of long-range magnetic order at the two-dimensional (2D) limit, which does not follow the Mermin-Wagner theorem [5] due to finite magnetic anisotropies, has stimulated an intense exploration of the growing variety of magnetic 2D materials [6, 7]. Similar to synthetic antiferromagnets (SyAF) [8], vdW AFM’s such as CrCl3 [9] and CrPS4 [10] have a weak interlayer exchange coupling, which brings the typical terahertz (THz) dynamics of atomistic AFM’s down to the gigahertz (GHz) regime. This characteristic makes them an ideal platform for studying a variety of spin dynamics phenomena using modern microwave techniques.
@@
The interaction between magnons and other quasi-particles has garnered interest due to its fundamental importance in understanding complex quantum material interactions and its potential application in hybrid magnonics [11, 12]. Through leveraging the unique properties of magnons, these systems have shown significant promise in quantum information processing, storage, and sensing [13, 14]. Due to the inherently weak dipolar interaction between magnons and photons, it is difficult to achieve ultrastrong magnon-photon coupling [15, 16]. The normalised coupling strength, defined as the ratio of the gap size to the bare excitation frequency, is typically used to quantify interaction regimes. When the normalised coupling strength increases above 0.1, the ultrastrong coupling regime is realized [17, 18, 19]. For light-matter systems in this regime, the rotating wave approximation no longer holds and the gap size is no longer a good representation of the coupling strength. To account for this more accurate models are adopted, such as the quantum Rabi model and Hopfield model [20]. In ultrastrong coupling systems, the interaction Hamiltonian consists of both the co-rotating and counter-rotating terms. This permits exotic physical phenomena, such as nontrivial ground states and superradiant phase transitions [21, 22, 23]. Recently the interaction between magnons themselves has emerged as a new avenue for investigating coupling phenomena [24, 9, 25, 26]. Unlike magnon-photon coupling, which are coupled through the relatively weak dipolar interaction, the intrinsically strong interaction between magnons results in an enhanced coupling strength, due to the spatial mode overlap through a shared host media [21].
@@
Quantifying the strength of magnon-magnon coupling remains a topic of considerable debate. The gap size is commonly used as a measure of magnon-magnon coupling strength, serving as a tool to define ultrastrong interactions in magnon-magnon systems [27, 28, 24, 21, 26, 29]. It has been shown that when intrinsic symmetry breaking induces the coupling, that an indirect gap opening occurs between the resonance modes with further increased coupling strength [30]. Additionally, a dc-field independent coupling strength parameter has been proposed for defining this coupling, which can be derived from material parameters [30].
In addition to magnonics, the conversion of charge to spin currents is a crucial cornerstone of spintronics particularly for artificial intelligence and high-performance computing applications requiring high-speed on-chip memory [31, 32, 33]. Spin Hall Magnetoresistance (SMR) serves as a powerful tool for probing magnetisation states and spin transport phenomena in magnetic heterostructures. SMR in ferromagnetic/non-magnetic structures relies on the relative orientations of magnetisation and accumulated spins at the interface [34]. Although recent investigations into SMR in vdW materials have emerged, there are still few comprehensive studies [35, 36, 37]. The presence of electrical readout of magnetisation presents this system as a promising platform for future spin dynamics studies on thin flakes, for example using Spin-Torque Ferromagnetic Resonance (ST-FMR), which is highly desirable for exploration of the 2D nature of magnetism in vdW flakes [38].
@@
In this work, we demonstrate ultrastrong magnon-magnon coupling in the vdW AFM material CrPS4 in the GHz regime, achieving a normalised coupling strength of 0.89. We show how the orthorhombic magnetic anisotropy lowers the symmetry of the magnetic system, thereby inducing magnon-magnon coupling between the two resonance modes. Furthermore, we investigate the temperature dependence of this coupling and report the magnetic parameters as a function of temperature in the AFM regime. We demonstrate the tunability of coupling strength with temperature through controlling magnetic parameters, such as saturation magnetisation and the magnetic anisotropy. We compare coupling strength values derived from gap size with those calculated directly from magnetic properties. This comparison shows that gap size as a measure of coupling strength is limited when intrinsic symmetry breaking drives magnon mode coupling. Additionally, we show that it is possible to tune this coupling strength the deep-strong regime with normalised coupling strengths greater than 1.
Moreover, we observe distinct spin transport effects on the magnetisation in a vdW AFM exfoliated flake at cryogenic temperatures using SMR. We fabricate a Pt/CrPS4/h-BN Hall bar device and measure the SMR as a function of magnetic field and temperature. This highlights the ability to electrically detect the magnetic state on exfoliated flakes within this system.
@@
Figure 1: Structure and magnetism in $\mathbf{CrPS}_{4}$. a) Schematic illustration of crystal and magnetic structures of $\mathrm{CrPS}_{4}$. The blue and red arrows indicate the orientation of magnetic moments. b) X-ray diffraction of a single bulk CrPS4 crystal. (c) Zero-field-cooled magnetisation ($M$) of a $\mathrm{CrPS}_{4}$ single crystal as a function of temperature ($T$), with a 10 mT magnetic field applied parallel to the crystallographic c-axis. The red line indicates $T_{\mathrm{N}}$. The inset illustrates the behavior under a magnetic field applied perpendicular to the c-axis. (d) Field-dependent normalised magnetisation $\left(M/\left(M_{S(5K)}\right)\right.)$ at various temperatures with the magnetic field $\left(\mu_{0}H\right)$ oriented parallel to the c-axis. The inset provides a zoomed-in image around spin-flip transition $\left(H_{\mathrm{sf}}\right)$. (e) Field-dependent normalised magnetisation at various temperatures with the field-oriented perpendicular to the $c$-axis. (f) Saturation magnetisation $\left(M_{\mathrm{s}}\right)$ and spin-flop transition field $\left(H_{\mathrm{sf}}\right)$ as a function of temperature ($T$).
@@
CrPS4 is characterized by its semiconducting properties and an optically measured band gap of approximately 1.4 eV [39, 40]. The crystal structure exhibits non-centrosymmetricity with monoclinic anisotropy, belonging to space group C2, with crystal axes of lengths $a=10.871$, $b=7.254$, $c=6.140$, and $\beta=91.88^{\circ}$ [41]. The crystal structure is illustrated in Fig. 1(a). In Fig. 1(b), we show X-ray diffraction of a single bulk CrPS4 crystal. The position of 00x peaks is in agreement with A. Louisy et al.[40] confirming the successful growth of high quality CrPS4.
@@
We initially focus on determining the magnetic properties of the bulk single crystal sample, we present the magnetometry results obtained from our crystal in two field orientations, $H//c$ and $H//ab$, as functions of the applied magnetic field, $H$, and temperature, $T$. In Fig. 1(c), we depict the magnetisation, $M$, versus temperature under a small applied field (10 mT) perpendicular to the vdW plane. A sharp peak at 38 K followed by a steep decline is observed, indicative of the Néel transition temperature, $T_{\mathrm{N}}$. This observed $T_{\mathrm{N}}$ of 38 K agrees well with literature-reported values [10]. Fig. 1(d) and 1(e) depict $M$ versus $H$ for the $H//c$ and $H//ab$ orientations, respectively, at temperatures ranging from 5 K to 50 K. For temperatures above 10 K, a prominent non-saturating background is observed. From these we extract the saturation magnetisation as a function of temperature. In the $H//c$ orientation, a characteristic spin-flop transition is evident due to the easy axis anisotropy of the material. The spin-flop transition is indicated by the plateau in the moment until reaching a critical field, $H_{\mathrm{sf}}$. Beyond $H_{\mathrm{sf}}$, a sharp increase in $M$ is observed as the sublattices become canted towards the direction $H$. With increasing $H$, a linear rise in $M$ is measured until saturation, as the angle between the sublattices and $H$ decreases, ultimately saturating in the forced FM phase around 8 T at 5 K. The extracted values of $M_{\mathrm{s}}$ and $H_{\mathrm{sf}}$ are plotted as functions of temperature in Fig. 1(f), where both values decrease with increasing temperature.
Spin dynamics in a vdW AFM
@@
We now focus on ferromagnetic resonance (FMR) measurements conducted below the $T_{\mathrm{N}}$ for two orientations: $H\perp c$ with $\theta\approx 90^{\circ}$ and $45^{\circ}$, where $\theta$ is the angle between the $b$ axis and $H$. The crystal axis of the bulk crystal is determined visually, considering the straight edge of the crystal as the crystal growth aligns with the preferred $b$ axis [10]. This method of determination differs from the preferential edges observed during exfoliation, which typically exhibit an angle of $33.75^{\circ}$ away from the $a$ axis [42, 39]. The experimental setup is depicted in Fig. 2(c).
Tunable ultrastrong magnon–magnon coupling
@@
Figure 2: Spin dynamics in $\mathbf{CrPS}_{4}$ under an in-plane magnetic field. (a-b) Schematic illustration of the magnetic moments’ precession in the acoustic mode (in-phase precession) (a) and the optical mode (out-of-phase precession) (b). (c) Schematic of the configuration between coplanar waveguide (CPW), $\mathrm{CrPS}_{4}$ crystal, and external in-plane magnetic field ($H$). (d) Typical ferromagnetic resonance (FMR) spectra (open circles) obtained at 15 K and 15 GHz and their derivated Lorentzian fitting (green line) by using Eq. S1 (see SI). (e-f) Frequency ($f$) dependent FMR results obtained at 15 K, with the magnetic field applied ($\mu_{0}H$) along the $a$-axis (e) and at $\theta=45^{\circ}$ (f). The dots represent experimental data, and the fitting using Eq. 4. The fitted resonance modes are denoted as BAM, BOM, HFHM, and LFHM, representing the bare acoustic mode, bare optical mode, high-frequency hybrid mode, and low-frequency hybrid mode, respectively. (g) $H_{\mathrm{A1}}$ and $H_{\mathrm{A2}}$, as functions of temperature. (h) Normalised magnon-magnon coupling strength, $\tilde{\Delta}$ and $\eta$, as functions of temperature at $\theta=45^{\circ}$. (i) Field dependence of ${\Delta}/2\pi$ compared to the $\delta f(H)$ at 15K, where $\delta f_{min}$ is the half of the minimum gap size.
@@
Below, we present a detailed study of the temperature dependence of the resonances measured in the in-plane orientation with respect to the crystal axis, $b$, at angles $\theta=90^{\circ}$ and $\theta\approx 45^{\circ}$. Measurements in the out-of-plane orientation are included in the SI (Fig. S5).
@@
To model the spin dynamics in an AFM system, we can employ a macrospin approach using the Landau-Liftshitz (LL) equation for two ferromagnetic sublattices. This approach results in coupled equations, allowing the calculation of frequency versus field spectra. Starting with Kittel’s free energy equation for orthorhombic anisotropy and neglecting the terms dependent upon the relative positions of the sublattice moments [43]:
@@
where $F_{\mathrm{A}}$ is the anisotropy component free energy, $K_{1}$ and $K_{2}$ are the anisotropy constants related to the $a$ and $b$ axis, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$ and $\beta_{2}$ are the cosines of the unit magnetisation for the two sublattices. The LL equation can then be written as follows:
@@
The magnon-magnon coupling can be described as an 2x2 eigenvalue problem as follows:
@@
where $\Delta$ denotes the magnon-magnon coupling term and $\omega_{\text{BAM}}$ and $\omega_{\text{BOM}}$ are the bare acoustic and optical mode frequencies respectively. These precession modes are depicted in Fig. 2(a) and 2(b) respectively.
@@
Solutions to this reveal that when $\theta\neq 0\degree,90\degree$ an anticrossing gap opens up at the intersection between the two modes through mode hybridisation. The anti-crossing gap is the signature of such hybridisation and the size of the gap of the gap between the hybridised modes is defined here as $2\delta f$ which we take at zero field as this is the point where the uncoupled modes would intersect at $\theta=45\degree$ Eq. 3 can then be used to fit experimental data as described in the following section. Solving for the real part of the above equation for $\theta=0^{\circ}$, we obtain the eigenfrequency of the bare acoustic and optical modes, BAM and BOM respectively, given by:
[…]
In Fig. 2(e), when $\theta=90^{\circ}$, we observe that only one mode can be measured, which corresponds well with the BAM described in Eq. 4. No BOM is detected because, within this field window, the mode’s frequency remains almost constant and thus cannot be detected by the field sweep measurements conducted. This observation confirms the absence of a symmetry-breaking condition in this orientation.
@@
In Fig. 2(f), we plot the spectra obtained at 15 K in the in-plane orientation with $\theta\approx 45^{\degree}$. Here, we clearly observe magnon-magnon coupling, as evidenced by the opening of the anticrossing gap, which is comparable to the bare magnon frequencies and hence in the ultrastrong coupling regime. These two hybridised modes can be well-fitted with the computation of Eq. 3. By fitting the coupled modes, we can extract the magnetic parameters, $H_{\mathrm{A1}}$, and $H_{\mathrm{A2}}$. Furthermore, we calculate the coupling strength as half the gap size, $\delta f$, of the coupled modes at the field at which the two uncoupled modes would intersect ( 0 T when $\theta=45\degree$).
The values obtained for the two in-plane anisotropy fields $H_{\mathrm{A1}}$ and $H_{\mathrm{A2}}$, are presented in Fig. 2(g). Throughout all temperatures, the value of $H_{\mathrm{A1}}$ remains larger than $H_{\mathrm{A2}}$. Both anisotropy fields follow a similar decreasing function of temperature, with a deviation observed in $H_{\mathrm{A2}}$ at 15 K and 20 K. It is typical to use the normalised coupling strength, $\eta$, by comparing the anticrossing gap size to the uncoupled excitation frequency. $\eta$ is given by the following:
$\eta=\frac{\delta f}{f_{\mathrm{r}}}$ (5)
@@
where $f_{\mathrm{r}}$ represents the bare excitation frequency at the crossing.
@@
We now define the intrinsic normalised coupling strength as follows from coupling strength defined in Eq. 3:
$\tilde{\Delta}=\frac{\Delta}{2\pi f_{\mathrm{r}}}$ (6)
@@
We then compare the normalised magnon-magnon coupling strengths as a function of temperature obtained through the two methods, $\eta$ and $\tilde{\Delta}$, as shown in Fig. 2(h). We observe a significant difference between the two methods for all temperatures, with $\tilde{\Delta}$ more than twice the value of $\eta$.

### v4

# Tunable Ultra-Strong Magnon-Magnon Coupling Approaching the Deep-Strong Regime in a van der Waals Antiferromagnet

@@
Tunable Ultra-Strong Magnon-Magnon Coupling Approaching the Deep-Strong Regime in a van der Waals Antiferromagnet
C. W. F. Freeman1,2,∗, H. Youel1, A.K. Budniak3,4, Z. Xue1, H. De Libero5, T. Thomson5,
@@
Antiferromagnetic (AFM) magnons in van der Waals (vdW) materials offer substantial potential for applications in magnonics and spintronics. In this study, we demonstrate ultra-strong magnon-magnon coupling in the GHz regime within a vdW AFM, achieving a maximum coupling rate of 0.91. Our investigation shows the tunability of coupling strength through temperature-dependent magnetic anisotropies. We compare coupling strength values derived from the gap size from the measured spectrum with those calculated directly through the coupling parameter and show that the gap size as a measure of coupling strength is limited for the ultra-strong coupling regime. Additionally, analytical calculations show the possibility to reach the deep-strong coupling regime by engineering the magnetic anisotropy. These findings highlight the potential of vdW AFMs as a model case to study magnetisation dynamics in low-symmetry magnetic materials.
@@
KEYWORDS: spintronics, hybrid magnonics, 2D materials, antiferromagnets, spin dynamics
@@
Within the expansive range of materials explored for spintronic applications, antiferromagnets (AFMs) have emerged as compelling candidates owing to their unique properties and potential to address challenges associated with traditional ferromagnetic (FM) devices [1, 2]. Characterised by antiparallel alignment of neighboring magnetic moments, net zero magnetisation and atomistic exchange coupling strength, AFMs offer inherent advantages such as insensitivity to external magnetic fields and ultra-fast spin dynamics [3]. However, the intricate nature of spin dynamics in AFMs presents both opportunities and challenges, underscoring the need for further investigation to fully harness their potential in spintronics.
@@
Layered van der Waals (vdW) materials can be mechanically exfoliated down to monolayers relatively easily under controlled experimental conditions [4]. The recent discovery of long-range magnetic order at the two-dimensional (2D) limit, which does not follow the Mermin-Wagner theorem [5] due to finite magnetic anisotropies, has stimulated an intense exploration of the growing variety of magnetic 2D materials [6, 7]. Similar to synthetic antiferromagnets (SyAF) [8], vdW AFM’s such as CrCl3 [9] and CrPS4 [10] have a weak interlayer exchange coupling, which brings the typical terahertz (THz) dynamics of atomistic AFM’s down to the gigahertz (GHz) regime. This characteristic makes them an ideal platform for studying a variety of spin dynamics phenomena using modern microwave techniques.
@@
The interaction between magnons and other quasiparticles has garnered interest due to its fundamental importance in understanding complex quantum material interactions and its potential application in hybrid magnonics [11, 12]. Using the unique properties of magnons, these systems have shown significant promise in quantum information processing, storage, and sensing [13, 14]. Due to the inherently weak dipolar interaction between magnons and photons, it is often difficult to achieve ultra-strong magnon-photon coupling without extensive engineering and optimisation [15, 16]. The coupling rate, typically defined as the ratio of the gap size to the bare excitation frequency, is used to quantify the rate of information transfer between coupled modes and defines the interaction regimes. When the coupling rate increases above 0.1, the ultra-strong coupling regime is realised [17, 18, 19]. For light-matter systems in this regime, the rotating wave approximation does not hold and the gap size is no longer a good representation of the coupling strength. To account for this more accurate models are adopted, such as the quantum Rabi model and Hopfield model [20]. In ultra-strong coupling systems, the interaction Hamiltonian consists of both the co-rotating and counter-rotating terms. This permits exotic physical phenomena, such as nontrivial ground states and superradiant phase transitions [21, 22, 23]. Recently the interaction between magnons themselves has emerged as a new avenue for investigating coupling phenomena [9, 24, 25, 26]. Unlike magnon-photon coupling, a strong magnon-magnon coupling strength is in general expected due to their large spatial mode overlap [21].
@@
Quantifying the strength of magnon-magnon coupling remains a topic of considerable debate. The gap size is commonly used as a measure of magnon-magnon coupling strength, serving as a tool to define ultra-strong interactions in magnon-magnon systems [21, 26, 27, 28, 30]. Recently, it has been shown that when intrinsic symmetry breaking induces the coupling, that an indirect gap opening occurs between the resonance modes with increased coupling strength than through extrinsic means [31]. Additionally, a dc-field independent coupling strength parameter has been proposed for defining this coupling, which can be derived from material parameters in the case of SyAF with asymmetric magnetisation between the layers [31].
@@
In this work, we demonstrate ultra-strong magnon-magnon coupling in the vdW AFM material CrPS4 in the GHz regime, achieving a coupling rate of 0.91. We show how the orthorhombic magnetic anisotropy lowers the symmetry of the magnetic system, thereby inducing magnon-magnon coupling between the two resonance modes. Furthermore, we investigate the temperature dependence of this coupling and report the magnetic parameters as a function of temperature. We demonstrate the tunability of coupling strength with temperature through controlling magnetic parameters, such as saturation magnetisation and the magnetic anisotropy. We compare coupling rate values derived from gap size with those calculated directly from the intrinsic coupling parameter and discuss the difference between the two. Our model predicts a viable strategy to reach the deep-strong coupling regime by changing magnetic anisotropies.
@@
Figure 1: Structure and magnetism in $\mathbf{CrPS}_{4}$. a) Schematic illustration of crystal and magnetic structures of $\mathrm{CrPS}_{4}$. The blue and red arrows indicate the orientation of magnetic moments. b) X-ray diffraction of a single bulk CrPS4 crystal. (c) Zero-field-cooled magnetisation ($M$) of a $\mathrm{CrPS}_{4}$ single crystal as a function of temperature ($T$), with a 10 mT magnetic field applied parallel to the crystallographic c-axis. The red line indicates $T_{\mathrm{N}}$. The inset illustrates the behavior under a magnetic field applied perpendicular to the c-axis. (d) Field-dependent normalised magnetisation $\left(M/\left(M_{S(5K)}\right)\right)$ at various temperatures with the magnetic field $\left(\mu_{0}H\right)$ oriented parallel to the c-axis. The inset provides a zoomed-in image around spin-flop transition $\left(H_{\mathrm{sf}}\right)$. (e) Field-dependent normalised magnetisation at various temperatures with the field-oriented perpendicular to the $c$-axis. (f) Saturation magnetisation $\left(M_{\mathrm{s}}\right)$ and spin-flop transition field $\left(H_{\mathrm{sf}}\right)$ as a function of temperature ($T$).
@@
CrPS4 is characterised by its semiconducting properties and an optically measured band gap of approximately 1.4 eV [32, 33]. The crystal structure exhibits non-centrosymmetricity with monoclinic anisotropy, belonging to space group C2, with crystal axes of lengths $a=10.871$, $b=7.254$, $c=6.140$, and $\beta=91.88^{\circ}$ [34]. The crystal structure is illustrated in Fig. 1(a). In Fig. 1(b), we show X-ray diffraction of a single bulk CrPS4 crystal. The position of 00x peaks is in agreement with work by Louisy et al. [33] confirming the successful growth of high-quality CrPS4.
@@
We present the magnetometry results obtained from our bulk crystal in two field orientations, $H//c$ and $H//ab$, as functions of the applied magnetic field, $H$, and temperature, $T$. In Fig. 1(c), we depict the magnetisation, $M$, versus $T$ under a small applied field (10 mT) for $H//c$. A sharp peak at 38 K is observed followed by a steep decline, indicative of the Néel transition temperature, $T_{\mathrm{N}}$, consistent with that in the literature [10]. Figures 1(d) and 1(e) plot $M$ versus $H$ for the $H//c$ and $H//ab$, respectively, at $T$ ranging from 5 K to 50 K. For $T$ above 10 K, a prominent non-saturating background is observed. From these, we extract the saturation magnetisation as a function of $T$. In the $H//c$ orientation, a characteristic spin-flop transition is evident due to the out-of-plane magnetic easy axis of the material [10]. The spin-flop transition is indicated by the plateau in the moment until reaching a critical field, $H_{\mathrm{sf}}$. Beyond $H_{\mathrm{sf}}$, a sharp increase in $M$ is observed as the sublattices become canted towards the direction $H$. With increasing $H$, a linear rise in $M$ is measured until saturation, as the angle between the sublattices and $H$ decreases, ultimately saturating in the forced FM phase around 8 T at 5 K. The extracted values of $M_{\mathrm{s}}$ and $H_{\mathrm{sf}}$ are plotted as functions of temperature in Fig. 1(f), where both values decrease with increasing temperature.
Tunable Ultra-Strong Magnon–Magnon Coupling Approaching the Deep-Strong Regime
@@
We focus on ferromagnetic resonance (FMR) measurements conducted below the $T_{\mathrm{N}}$ for two orientations: $H\perp c$ with $\theta\approx 90^{\circ}$ and $45^{\circ}$, where $\theta$ is the angle between the $b$ axis and $H$. The crystal axis of the bulk crystal is determined visually, considering the straight edge of the crystal as the crystal growth aligns with the preferred $b$ axis [10]. This method of determination differs from the preferential edges observed during exfoliation, which typically exhibit an angle of $33.75^{\circ}$ away from the $a$ axis [32, 35].
@@
Figure 2: Spin dynamics in $\mathbf{CrPS}_{4}$ under an in-plane magnetic field. (a-b) Schematic illustration of the magnetic moments’ precession in the acoustic mode (in-phase precession) (a) and the optical mode (out-of-phase precession) (b). (c) Schematic of the configuration between coplanar waveguide (CPW), $\mathrm{CrPS}_{4}$ crystal, and external in-plane magnetic field ($H$). The angle between $H$ and the $b$ axis is shown by $\theta$. (d) Typical ferromagnetic resonance (FMR) spectra (red line) obtained at 15 K and 200 mT and their derivated Lorentzian fitting (black line) by using Eq. S9 (see SI). (e-f) 2D maps of the FMR spectra obtained at 15 K, with the magnetic field applied ($\mu_{0}H$) at $\theta=90^{\circ}$ (e) and $\theta=45^{\circ}$ (f). The fitted resonance modes are denoted as BAM, BOM, HFHM, and LFHM, representing the bare acoustic mode, bare optical mode, high-frequency hybrid mode, and low-frequency hybrid mode, respectively. The gap size $g/\pi$ is depicted in (f) at 0 T where the uncoupled modes would intersect. (g-h) Simulated FMR spectra obtained with the parameters extracted from experimental data with $\mu_{0}H$ at $\theta=90^{\circ}$ (g) and $\theta=45^{\circ}$ (h).
Figure 3: Temperature dependence of magnetic parameters and coupling (a-b) Plot showing the temperature dependence of $\mu_{0}H_{A1}$, $\mu_{0}H_{A2}$ (a) and $\mu_{0}H_{E}$ (b). (c) Temperature dependence of the intrinsic coupling strength, $\Delta_{0}/2\pi$, coupling strength, $g/2\pi$, and the line width, $\kappa/2\pi$. A significant increase between the two methods of calculating coupling strength is observed. (d) Temperature dependence of the coupling rate determined through the two methods, gap size $\eta$ and intrinsic coupling $\tilde{\Delta}$. The intrinsic coupling rate is seen to approach the deep-strong coupling regime with temperature. SC, USC, and DSC represent strong coupling, ultra-strong coupling, and deep-strong coupling regimes, respectively.
@@
We now present a detailed study of the temperature dependence of the resonances measured in the in-plane orientation with respect to the crystal axis, $b$, at angles $\theta\approx 90^{\circ}$ and $45^{\circ}$. Measurements in the out-of-plane orientation are included in the Supplementary Infomation (SI) (Figs. S1 and S2).
@@
To model the spin dynamics, we employ the coupled Landau-Liftshitz-Gilbert (LLG) equation for two macrospin sublattices coupled through an interlayer exchange interaction, with zero damping for simplicity. Starting with Keffer’s free energy equation for orthorhombic anisotropy and neglecting the terms dependent upon the relative positions of the sublattice moments [36]:
@@
where $F_{\mathrm{A}}$ is the anisotropy component free energy, $K_{1}$ and $K_{2}$ are the anisotropy constants related to the $a$ and $b$ axis, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$ and $\beta_{2}$ are the cosines of the unit magnetisation for the two sublattices. The LLG equation can then be written as follows:
@@
After Fourier transformation, the LLG equation can be treated as an 2$\times$2 eigenvalue problem with the characteristic equation of:
@@
where $\Delta$ denotes the intrinsic magnon-magnon coupling term and $\omega_{\text{BAM}}$ and $\omega_{\text{BOM}}$ are the bare acoustic and optical mode frequencies respectively. These precession modes are depicted in Fig. 2(a) and 2(b) respectively.
@@
Solutions to this reveal that when $\theta\neq 0^{\circ}$ an anti-crossing gap opens up around the intersection between the two modes through mode hybridisation. The anti-crossing gap is the signature of such hybridisation and the gap size between the hybridised modes is defined here as $g/\pi$ which we take at zero field as this is the point where the uncoupled modes would intersect at $\theta=45^{\circ}$ Equation 3 can then be used to fit experimental data as described in the following section. Solving Eq. 3 for $\theta=0^{\circ}$, we can obtain the frequency versus field dependence of the bare acoustic and optical modes, BAM and BOM respectively, given by Eq. S8 in the SI.
The experimental setup is depicted in Fig. 2(c). A typical frequency swept spectra is shown in Fig. 2(d) which can be fitted with two differentiated Lorentzian peaks (see SI, Eq. S9), from which the resonance frequency can be extracted. In Fig. 2(e), we plot a 2D map of the spectra obtained when $\theta\approx 90^{\circ}$ at 15 K. We observe two modes that can be fitted by BAM and BOM described in Eq. 3. Furthermore, simulations at $\theta=0^{\circ}$ shown in Fig. S3(b) in the SI demonstrate the intersection of the two modes along this high symmetry axis. This observation confirms the absence of a symmetry-breaking condition in these orientations.
@@
In Fig. 2(f), we plot the 2D map of the spectra obtained at 15 K in the in-plane orientation with $\theta\approx 45^{\circ}$. Here, we clearly observe magnon-magnon coupling, as evidenced by the opening of the anticrossing gap, which is comparable to the bare magnon frequencies and hence in the ultra-strong coupling regime. These two hybridised modes can be well-fitted with the computation of Eq. 3.
Two other modes are also observed in the 2D map. The mode that crosses the LFHM at approximately 1.5 T in Fig. 2(f) is identified as a harmonic resonance of the HFHM mode and appears at half the resonant frequency of the HFHM. Additionally, double, quarter, and eighth harmonics are observed, which are fully depicted in Fig. S4 of the SI. The other mode, a linear mode reaching 40 GHz at approximately 1.2 T, is explained by background absorption originating from the waveguide itself. Reference measurements are provided to confirm this (see SI, Fig. S5).
In Figs. 2(g)-(h), we present micromagnetic simulations carried out in Mumax3, for $\theta=90^{\circ}$ and $\theta=45^{\circ}$, respectively. The simulation is carried out using the magnetic parameters determined from fitting at $15$ K and shows good agreement with the experimental data as shown in Fig. 2(e-f).
By fitting the experimental resonant frequencies for the coupled modes, we can extract $H_{\mathrm{A1}}$ and $H_{\mathrm{A2}}$. Furthermore, we calculate the coupling strength as half the gap size, $g/2\pi$, of the coupled modes at the field at which the two uncoupled modes would intersect (0 T when $\theta=45^{\circ}$).
The values obtained for $H_{\mathrm{A1}}$ and $H_{\mathrm{A2}}$, are presented in Fig. 3(a). Throughout all temperatures, the value of $H_{\mathrm{A1}}$ remains larger than $H_{\mathrm{A2}}$. Both anisotropy fields follow a decreasing function with temperature as is typical of magneto-crystalline anisotropies [37]. In Fig. 3(b), we plot the values of $H_{E}$ versus $T$ which also decreases as a function of temperature as would be expected and agreeing with recent results using magnetic torque [38].
It is typical to use the coupling rate, $\eta$, by comparing the anticrossing gap size to the uncoupled excitation frequency, where $\eta$ is given by the following:
$\eta=\frac{g}{2\pi f_{\mathrm{r}}}$ (4)
@@
Here, $f_{\mathrm{r}}$ represents the bare excitation frequency at the crossing of the uncoupled BAM and BOM, for $\theta=45^{\circ}$ this occurs at 0 T.
@@
Similarly, we can now define the intrinsic coupling rate, $\tilde{\Delta}$, as follows from the intrinsic magnon-magnon coupling strength defined in Eq. 3:
$\tilde{\Delta}=\frac{\Delta_{0}}{2\pi f_{\mathrm{r}}}$ (5)
where $\Delta_{0}$ is the value of $\Delta$ at the point of intersection of the uncoupled modes.
In Fig. 3(c), we compare the values of $\Delta_{0}/2\pi$ and $g/2\pi$ to the linewidth, $\kappa/2\pi$, of the hybridised modes. The linewidth is taken at a field of 200 mT. No clear field or temperature dependence of the linewidth is observed due to large extrinsic broadening (see SI, Fig. S6) as is common in bulk vdW samples where inhomogeneity across the sample makes linewidth analysis difficult [39]. In both $\Delta_{0}/2\pi$ and $g/2\pi$, we show that the coupling strengths are significantly larger than $\kappa/2\pi$ meeting the criteria for strong coupling [22, 19]. Furthermore, we observe that $\Delta_{0}/2\pi$ is significantly larger than $g/2\pi$ over the the entire temperature range. This agrees with the findings of Li et al. [31] in asymmetric SyAFs, which suggest that the gap size should not be used to quantify the coupling strength in cases of intrinsic symmetry-breaking-induced magnon-magnon coupling.
Figure 4: Tunable coupling rate via magnetic anisotropy up to deep-strong coupling. (a) Field dependence of ${\Delta}/2\pi$ compared to the $g/2\pi$ at 15K, where $g_{min}/2\pi$ is the half of the minimum gap size. (b) Analytical values of $\eta$ and $\tilde{\Delta}$ as a function of $H_{A2}$, when $\mu_{0}H_{A2}$ is equal to $\mu_{0}M_{s}$ the deep strong regime is reached. (c) Micromagnetic simulation of the FMR spectra for the deep-strong coupling case, which shows good agreement with the analytical solutions. (d) 2D plot of $\tilde{\Delta}$ for a fixed value of $M_{\mathrm{s}}$ at 33 K as a function of $H_{\mathrm{A2}}$ and $H_{\mathrm{A1}}$, demonstrating the potential tunability of the coupling rate. The star indicates the position and value of $\tilde{\Delta}$ observed experimentally at 33 K.
@@
We now compare the coupling rates as a function of temperature obtained through the two methods, $\eta$ and $\tilde{\Delta}$, as shown in Fig. 3(d). We observe a significant difference between the two methods at all temperatures, with $\tilde{\Delta}$ being more than twice the value of $\eta$. Furthermore, $\tilde{\Delta}$ increases as a function of $T$, from a minimum of 0.55 at 5 K to a maximum of 0.91 at 33 K, approaching the deep-strong regime as the temperature approaches $T_{N}$. This temperature dependence arises from the interplay between the magnetic parameters as a function of temperature, as we demonstrate below.